This question seemed simple enough but....help is needed!
The Schwarzchild radius describes the critical value to which the radius of a massive body must be reduced for it to become a black hole.
R = 2GM / c^2
where
G = gravitional constant 6.7 x 10^-11
M = mass of the onject
C = speed of light 3 x 10^8
The sun has M = 2 x 10^30. What is the Schwarzchilds radius for the sun? The suns true radius is 700,000.
: This question seemed simple enough but....help is needed!
The Schwarzchild radius describes the critical value to which the radius of a massive body must be reduced for it to become a black hole.
R = 2GM / c^2
where
G = gravitional constant 6.7 x 10^-11
M = mass of the onject
C = speed of light 3 x 10^8
The sun has M = 2 x 10^30. What is the Schwarzchilds radius for the sun? The suns true radius is 700,000.

Answer by ankor@dixie-net.com(4668)

The Schwarzchild radius describes the critical value to which the radius of a massive body must be reduced for it to become a black hole.
R = 2GM / c^2
where
G = gravitional constant 6.7 x 10^-11
M = mass of the onject
C = speed of light 3 x 10^8
The sun has M = 2 x 10^30. What is the Schwarzchilds radius for the sun? The suns true radius is 700,000.
:
Wouldn't just be:
R = (2(6.7(10^-11))* 2(10^30))/(3(10^8))^2 = (2(6.7(10^-11))* 2(10^30))/(9(10^16)) = 26.8(10^((30-11-16)))/9 = 2.98(10^3) m
Which is about 2.98 km